(a) A straight conducting wire is placed perpendicularly to the earth

magnetic field at a location. The earth magnetic field at the location has a

magnitude of 0.450 G. Determine the current that must flow in the wire for a

9.50 cm length of the wire to experience a magnetic force of 1.50 N due to the

earth magnetic field.

[1 T = 104 G]


(b) A wire of length 1.0 m carries a current of 1.0 A. The wire is formed into a

circular coil and the circular coil has one turn only. Determine the magnitude of

the maximum torque which is developed on the coil if it is placed in magnetic

field of 2.0 T.


1
Expert's answer
2022-02-21T14:55:29-0500

Now we know that

Force

"F=ilB"

Put value

"i=\frac{F}{lB}"


"I=\frac{1.50}{9.50\times10^{-2}\times0.450\times10^{-4}}=3.5\times10^{5}Amp"


Part (b)

Torque

"\tau=\frac{L^2IB}{4\pi}"

Put value

"\tau=\frac{1^2\times1^2\times2}{4\times3.14}=0.16Am^2"